### v2

# Reduction-induced Variation of Partial Von Neumann Entropy

Abstract
The organization and structure of bipartite mixed-state quantum entanglement (QE) is more complicated and less well understood than that of bipartite pure-state QE. Bipartite mixed-state QEs and their measures play a crucial role in both theory and practical applications. Existing measures comprise two classes: one class involves quantifying the minimum QE and reflect the inherently complex nature of their computation, while the other is only applicable to highly limited-dimensional quantum systems. In this context, we propose a method termed Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE) to quantify QE in any bipartite states especially for any bipartite mixed-states, which is essentially an extension of Partial Von Neumann Entropy for measuring QE from bipartite pure states to any bipartite states. This method exhibits minimal computational complexity and broad applicability. Its intuitive and clear physical representation, combined with easy computation and wide applicability, facilitates exploring its potential applications. Furthermore, we present examples to demonstrate the superiorities of this method in identifying bipartite QE by comparing with other existing bipartite mixed-state QE measures through both their physical implications and mathematical structures.
Keywords:Quantum Entanglement (QE); Bipartite Mixed-state QE; Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE)
pacs:
03.65.Ud; 75.10.Pq; 05.30.-d
I Introduction
In 1935, “Spooky Action at a Distance”described as a strange phenomenon of quantum mechanics calls into question the completeness of the quantum theory by the famous physicists Einstein, Podolsky and Rosen (EPR) epr . And then in the same year, the concept of quantum entanglement (QE) was first proposed clearly by Schrödinger sd , through rigorous mathematical formalism, he demonstrated that particles in such entangled states cannot be individually described but must be treated as an integral composite system. Subsequently, in 1964, Bell realized that QE can produce experimentally verifiable deviations from classical physics in quantum mechanics bell . However, even to this day, its mystery still challenges our understanding of the laws of nature.
The type or degree of QEs are often directly related to their revolutionary applications including quantum cryptography kk ; ng , quantum teleportation ch ; rc ; nr , quantum superdense coding nr ; sj , measurement-based quantum computation rr ; fa and quantum enhanced sensing tc ; rcp ; lp . Therefore, a deeper understanding and accurate metric of QE have become important prerequisites for the development of quantum information science and quantum technology.
Nonseparable bipartite systems exhibit QE epr ; sd , which can be quantified using various measures. These include the Partial Von Neumann Entropy ch1 ; ch2 , applicable only to bipartite pure states, and Concurrence, which is confined to $2\times 2$ quantum systems wk ; au ; ka ; mf ; xjw , and Negativity, which is applicable only to $2\times 2$ and $2\times 3$ quantum systems K.Z ; gv . Additionally, Entanglement of Formation ch1 ; ch2 and Relative Entropy of Entanglement vv1 ; vv2 for bipartite quantum states, involve quantifying the minimum QE, and their assessments generally entail considerable computational burdens, especially for bipartite mixed-state QE. In general, QEs in pure states are more intuitive and easier to comprehend, while those in mixed states are more complicated and less well understood than those of pure-state QEs. Quantifying QEs in bipartite mixed states is more complex and challenging due to the contributions from classical probability distributions, making the definition and computation of QEs more intricate.
Mixed states are more reflective of real-world scenarios, as factors such as environmental conditions, quantum measurements, and thermal influences can lead quantum systems to exist in mixed states. Consequently, the measures of QEs in mixed states are often more directly applicable to complex systems, aligning better with experimental conditions and practical applications. Hence, in practical research and applications, bipartite mixed-state QE and its measures hold significant practical importance, besides the theoretical significance of their underlying structure of quantum mechanics.
In this context, our research aims to explore new metrics for bipartite mixed-state QE. We propose a method termed Reduction-induced Variation of Partial Von Neumann Entropy to quantify QE in any bipartite states, which is essentially an extension of Partial Von Neumann Entropy for measuring QE from bipartite pure states to any bipartite states. The computational complexity associated with this method is considerably low, making it highly efficient. Moreover, it has a wide range of applicability. The intuitive and clear physical representation, combined with simple computation and wide range of applications, facilitates the exploration of its specific potential applications. Subsequently, we provide examples to demonstrate the superiorities of this method in bipartite QE identification through a comparison with other established bipartite mixed-state QE measures including Partial Von Neumann Entropy, Concurrence, Negativity, Entanglement of Formation, and Relative Entropy of Entanglement, considering both their physical implications and mathematical structures.
II existing five kinds of Bipartite QE measures
II.1 Partial Von Neumann Entropy for Bipartite Pure-state QE
For a pure bipartite state $\mid\Psi\rangle$, its density matrix $\rho=\mid\Psi\rangle\langle\Psi\mid$, a well-known and commonly used bipartite pure-state QE measure, the Partial Von Neumann Entropy (PVNE) ch1 ; ch2 is defined as
$\operatorname{PVNE}(\rho)=S(\rho_{i})=-\operatorname{Tr}\rho_{i}\log_{2}{\rho_{i}},$ (1)
where $\rho_{i}(i=1,2)$ represents its subsystem reduced density matrix. This bipartite QE measure, the Partial Von Neumann Entropy (PVNE) can effectively measure the degree of QE between bipartite pure states, while for mixed-state QE, it is ineffective.
II.2 Entanglement of Formation for Bipartite QE and Concurrence for two-qubit QE
For bipartite QE especially for bipartite mixed-state QE, an extensively acknowledged type of measure, Entanglement of Formation (EOF) ch1 ; ch2 was introduced and is defined as
$\operatorname{EOF}(\rho)=\min_{\{p_{j},\mid\Psi_{j}\rangle\}}\sum_{j}p_{j}\operatorname{PVNE}(\mid\Psi_{j}\rangle\langle\Psi_{j}\mid),$ (2)
over all possible ensemble realizations of the bipartite density matrix $\rho=\sum_{j}p_{j}\mid\Psi_{j}\rangle\langle\Psi_{j}\mid$, where $p_{j}>0$ and $\sum_{j}p_{j}=1$. Entanglement of Formation represents the minimum entanglement resources required to construct this state and reflects that the involving computation is inherently complex. While for bipartite pure-state QE, Entanglement of Formation simplifies to the Partial Von Neumann Entropy (PVNE)
$\operatorname{EOF}(\mid\Psi\rangle)=\operatorname{PVNE}(\mid\Psi\rangle)=S(\rho_{i})=-\operatorname{Tr}\rho_{i}\log_{2}{\rho_{i}}.$ (3)
However, for mixed-state QE, in general, this measure heavily relies on the decompositions of pure states, with no general algorithm existing to determine the minimum decomposition to date. To address this issue and improve the Entanglement of Formation behavior in mixed states, Reference wk discovered that for two-qubit mixed states, Entanglement of Formation reduces to
$\operatorname{EOF}(\rho)=H(z)=-z\log(z)-(1-z)\log(1-z),$ (4)
where the argument z is given by
$z=\frac{1+\sqrt{1-C^{2}}}{2},$ (5)
and $C(\rho)$ denoting Concurrence is defined as wk ; au ; ka ; mf ; xjw
$\operatorname{C}(\rho)=\max\{0,\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4}\},$ (6)
where $\{\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\}$ represent the square roots of the eigenvalues of the matrix $\rho\tilde{\rho}$ arranged in decreasing order. The matrix $\tilde{\rho}$ is defined as
$\tilde{\rho}=(\sigma_{y}\otimes\sigma_{y})\rho^{*}(\sigma_{y}\otimes\sigma_{y}),$ (7)
where $\sigma_{y}$ is the Pauli matrix and $\rho^{*}$ denotes the complex conjugate of $\rho$. It is worth noting that, for two-qubit states, in itself, Concurrence is also a widely accepted and comparably practical measure of QE.
II.3 Relative Entropy of Entanglement for Bipartite QE
For bipartite QE especially for bipartite mixed-state QE, another well-known and commonly used bipartite QE measure, Relative Entropy of Entanglement (REE) vv1 ; vv2 was introduced and is defined as
$\operatorname{REE}=S(\rho|\sigma)=\min_{\sigma\in\mathcal{S}}[-\operatorname{Tr}\rho\log_{2}{\sigma}-S(\rho)],$ (8)
where $\mathcal{S}$ represents a set of separable bipartite density matrix. On one hand, Relative Entropy of Entanglement measures the shortest ”distance” from a given quantum entangled state to the set of separable states, which is the geometric interpretation of this method and reflects that the involving computation is inherently complex. On the other hand, while for bipartite pure QE states, Relative Entropy of Entanglement (REE) simplifies to the Partial Von Neumann Entropy (PVNE)
$\operatorname{REE}$ $=S(\rho|\sigma)=\min_{\sigma\in\mathcal{S}}[-\operatorname{Tr}\rho\log_{2}{\sigma}]$ (9)
$=\operatorname{PVNE}(\rho)=S(\rho_{i})=-\operatorname{Tr}\rho_{i}\log_{2}{\rho_{i}}.$
II.4 Negativity for $2\times 2$ and $2\times 3$ quantum systems
For a disentangled state $\rho$, its partial transpose results in a positive or valid state $\rho^{T_{i}}$, the so-called well-known Peres condition for separable states A.P . For $2\times 2$ and $2\times 3$ quantum states, this condition holds true M.H . It reveals a fact that the partial transpose of a non-separable state $\rho^{T_{i}}$ has one negative eigenvalue $\lambda_{\min}$. Then for $2\times 2$ and $2\times 3$ quantum states, Negativity is defined as twice the absolute value of this negative eigenvalue K.Z ; gv . Mathematically, it is expressed as:
$\operatorname{N}(\rho)=2\max(0,-\lambda_{\min}).$ (10)
Summary: Due to the locality of the aforementioned five kinds of bipartite QE measures, symmetry about different partitions is generally observed only for symmetric states. The widely recognized measures for quantifying bipartite mixed-state QE, such as Entanglement of Formation and Relative Entropy of Entanglement, involve determining the minimum QE required for a given state, reflecting the inherent complexity of their computation. While Concurrence is applicable only to $2\times 2$ quantum systems and Negativity limited to $2\times 2$ and $2\times 3$ quantum systems. The organization and structure of bipartite mixed-state QE is more complicated and less well understood than that of pure-state QE. Quantifying QE in mixed states poses additional challenges due to the contributions from classical probability distributions, which complicate both the definition and computation of QE measures. In this context, our research aims to explore new measures for bipartite mixed-state QE in detail, as outlined below.
III Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE)
In any bipartite quantum system, all the information is contained within its wave function or density matrix. When a subsystem is obtained by partial tracing or averaging out the other subsystem freedom, some information is inevitably lost. Therefore, the concept of Reduction-induced Variation of Partial Von Neumann Entropy is presented to quantify bipartite QE, which holds certain importance and practical significance. To elaborate further, let’s consider a bipartite quantum system. In the bipartite density matrix $\rho$, the density matrix of the $i$-th $(i=1,2)$ party, denoted as the bipartite $i$-th party density matrix, only preserves the information of the $i$-th party from $\rho$. This particular density matrix is represented as $\rho(i)$ different from its reduced density matrix $\rho_{i}$. Reduction-induced Variation of Von Neumann Entropy (RIVPVNE) of one-sided density matrix is defined as follows:
$\operatorname{RIVPVNE}(\rho)=S(\rho_{i})-S(\rho(i)),$ (11)
which is essentially an extension of Partial Von Neumann Entropy for measuring QE from bipartite pure states to any bipartite states especially for any bipartite mixed states. This measure vanishes if and only if the state is separable, and on average, does not experience an increase under stochastic Local Operations and Classical Communications, thereby establishing itself as an entanglement monotone. Notably, the computational complexity associated with this method is minimal, making it highly efficient. Moreover, it has a wide range of applicability. The intuitive and clear physical picture, combined with simple computation and wide range of applications, facilitates the exploration of its specific potential applications. For any bipartite pure-state QE, this measure reduces to Partial Von Neumann Entropy
$\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=-\operatorname{Tr}\rho_{i}\log_{2}{\rho_{i}}.$ (12)
Due to the locality of this measure, in general only for symmetric states, it has symmetry about different partition. Subsequently, we provide examples to demonstrate the superiorities of this method in identifying QE through a comparison with other existing bipartite QE measures, such as Partial Von Neumann Entropy, Concurrence, Negativity, Entanglement of Formation and Relative Entropy of Entanglement, considering both their physical implications and mathematical structures, as detailed below.
IV Its Superiorities in comparison to existing bipartite QE measures
We initially consider the symmetric pure Bell-like states
$|\Psi\rangle=\cos\alpha|11\rangle\pm\sin\alpha|00\rangle$ (13)
and
$|\Phi\rangle=\cos\alpha|10\rangle\pm\sin\alpha|01\rangle$ (14)
with $0\leqslant\alpha\leqslant 2\pi$. The corresponding single-site reduced density matrix and the bipartite single-site density matrix are respectively given by
$\rho_{i}=\left[\begin{array}[]{cc}\cos^{2}\alpha&0\\ 0&\sin^{2}\alpha\\ \end{array}\right]$ (15)
and
$\rho(i)=\left[\begin{array}[]{cc}\cos^{2}\alpha&\cos\alpha\sin\alpha\\ \cos\alpha\sin\alpha&\sin^{2}\alpha\\ \end{array}\right].$ (16)
Consequently, the Reduction-induced Variation of Partial Von Neumann Entropy, as well as the Entanglement of Formation and Relative Entropy of Entanglement, simplifies to the Partial Von Neumann Entropy:
$\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)$
$=-(\cos^{2}\alpha\log_{2}{\cos^{2}\alpha}+\sin^{2}\alpha\log_{2}{\sin^{2}\alpha}).$ (17)
Additionally, Concurrence and Negativity can be expressed as
$\operatorname{C}(\rho)=\operatorname{N}(\rho)=|\sin{2\alpha|}.$ (18)
This highlights the interference properties and indicates a clear and intuitive relationship between QE and quantum coherence. The previously mentioned measures are illustrated in Fig.1. For the specified pure states, it is observed that when $\alpha=\frac{\pi}{4},\frac{3\pi}{4}$, all the measures satisfy: $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=1$; whereas for $\alpha=0,\frac{\pi}{2},\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=0$; and for the ranges $0<\alpha<\frac{\pi}{2}$ and $\frac{\pi}{2}<\alpha<\pi$, the relationships $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)<\operatorname{C}(\rho)=\operatorname{N}(\rho)$ hold. Clearly, for the specified pure two-qubit systems, the monotonicities of these measures are fully consistent. The Reduction-induced Variation of Partial Von Neumann Entropy, Entanglement of Formation and Relative Entropy of Entanglement reduce to the Partial Von Neumann Entropy, but exhibit slight differences from Concurrence and Negativity due to variations in their physical implications and mathematical structures.
Fig. 1: Reduction-induced Variation of Partial Von Neumann Entropy as well as other QE measures such as the Partial Von Neumann Entropy, Entanglement of Formation, Relative Entropy of Entanglement, Concurrence and Negativity varying with $\alpha$. When $\alpha=\frac{\pi}{4},\frac{3\pi}{4}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=1$; whereas for the values $\alpha=0,\frac{\pi}{2},\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=0$; and in the ranges $0<\alpha<\frac{\pi}{2}$ and $\frac{\pi}{2}<\alpha<\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)<\operatorname{C}(\rho)=\operatorname{N}(\rho)$.
We next consider the symmetric bipartite mixed states
$\rho=\frac{\sin^{2}\alpha}{2}\mid 10\pm 01\rangle\langle 10\pm 01|+\cos^{2}\alpha|00\rangle\langle 00\mid.$ (19)
The first term corresponds to a QE state with a probability of $\sin^{2}\alpha$, while the second term represents a non-entangled state with a probability of $\cos^{2}\alpha$. Consequently, the bipartite QE is less than $\sin^{2}\alpha$. The corresponding single-site reduced density matrix and bipartite single-site density matrix are
$\rho_{i}=\left[\begin{array}[]{cc}\frac{\sin^{2}\alpha}{2}&0\\ 0&\frac{\sin^{2}\alpha}{2}+\cos^{2}\alpha\\ \end{array}\right]$ (20)
and
$\rho(i)=\left[\begin{array}[]{cc}\frac{\sin^{2}\alpha}{2}&\frac{\sin^{2}\alpha}{2}\\ \frac{\sin^{2}\alpha}{2}&\frac{\sin^{2}\alpha}{2}+\cos^{2}\alpha\\ \end{array}\right].$ (21)
Next, the Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE) is computed as
$\operatorname{RIVPVNE}=S(\rho_{i})-S(\rho(i))$
$=-(\frac{3+\cos{2\alpha}}{4}\log_{2}{\frac{3+\cos{2\alpha}}{4}}+\frac{\sin^{2}\alpha}{2}\log_{2}{\frac{\sin^{2}\alpha}{2}})$
$+\frac{2-\sqrt{3+\cos{4\alpha}}}{4}\log_{2}{\frac{2-\sqrt{3+\cos{4\alpha}}}{4}}$
$+\frac{2+\sqrt{3+\cos{4\alpha}}}{4}\log_{2}{\frac{2+\sqrt{3+\cos{4\alpha}}}{4}},$ (22)
as well as other QE measures such as the Entanglement of Formation
$\operatorname{EOF}(\rho)=$ $-\frac{1-\sqrt{1-\sin^{4}\alpha}}{2}\log_{2}{\frac{1-\sqrt{1-\sin^{4}\alpha}}{2}}$ (23)
$-\frac{1+\sqrt{1-\sin^{4}\alpha}}{2}\log_{2}{\frac{1+\sqrt{1-\sin^{4}\alpha}}{2}},$
the Relative Entropy of Entanglement and Concurrence:
$\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\sin^{2}\alpha,$ (24)
which represents the upper limit of QE as derived from the mathematical structure of the system’s density matrix and reflect that they may not accurately reflect the degree of mixed-state QE, and the Negativity expressed as
$\operatorname{N}(\rho)=\frac{\sqrt{3+\cos{4\alpha}}-1-\cos{2\alpha}}{2}.$ (25)
The five aforementioned measures are illustrated in Fig.2. Notably, when $\alpha=\frac{\pi}{2}$, the following relations hold: $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=1$; whereas for $\alpha=0,\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0$. For values $\alpha=0.8,2.341$, the relationships $\operatorname{RIVPVNE}(\rho)=\operatorname{N}(\rho)=0.222676<\operatorname{EOF}(\rho)=0.371412<\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0.515241$ hold; furthermore, in the ranges $0.80<\alpha<$ 0.8 and $2.341<\alpha<\pi$, we find $\operatorname{N}(\rho)<\operatorname{RIVPVNE}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$; in contrast, within the range $2.3410.8<\alpha<$ 2.341, the relations become $\operatorname{RIVPVNE}(\rho)<\operatorname{N}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$. Clearly, for the specified mixed two-qubit quantum systems, the monotonic behavior of these five measures is consistent. Specifically, the differences in the values of Reduction-induced Variation of Partial Von Neumann Entropy, Entanglement of Formation, Relative Entropy of Entanglement, Concurrence, and Negativity arise from their distinct physical implications and mathematical structures. Regarding the bipartite QE degree metric for the specified mixed states, it is concluded that the Reduction-induced Variation of Partial Von Neumann Entropy, Entanglement of Formation, and Negativity exhibit advantages over the Relative Entropy of Entanglement and Concurrence. When considering bipartite QE degree metric in mixed states, the computational complexity and range of applicability, the Reduction-induced Variation of Partial Von Neumann Entropy, is superior to Entanglement of Formation, Negativity, Relative Entropy of Entanglement and Concurrence.
Fig. 2: Reduction-induced Variation of Partial Von Neumann entropy as well as other QE measures such as Entanglement of Formation, Relative Entropy of Entanglement, Concurrence and Negativity varying with $\alpha$. When $\alpha=\frac{\pi}{2}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=1$; whereas for $\alpha=0,\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\operatorname{N}(\rho)=0$. For values $\alpha=0.8,2.341$, $\operatorname{RIVPVNE}(\rho)=\operatorname{N}(\rho)=0.222676<\operatorname{EOF}(\rho)=0.371412<\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0.515241$ ; furthermore, in the ranges $0.80<\alpha<$ 0.8 and $2.341<\alpha<\pi$, $\operatorname{N}(\rho)<\operatorname{RIVPVNE}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$; in contrast, within the range $2.3410.8<\alpha<$ 2.341, $\operatorname{RIVPVNE}(\rho)<\operatorname{N}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$.
V Conclusion
In this paper, we propose a method called Reduction-induced Variation of Partial Von Neumann Entropy to quantify QE in bipartite mixed states, which is essentially an extension of Partial Von Neumann Entropy for measuring QE from bipartite pure states to any bipartite states especially for any mixed states. This method exhibits minimal computational complexity, making it highly efficient. Moreover, it has a wide range of applicability. Its intuitive and clear physical representation, combined with simple computation and wide range of applications, indicates the ease of exploring its specific potential applications. By considering the degree of bipartite QE, along with the computational complexity and range of applicability, the Reduction-induced Variation of Partial Von Neumann Entropy proves to be a good measure for assessing bipartite mixed-state QE compared to established methods. In future studies, the conceptual framework of the Reduction-Induced Variation of Partial Von Neumann Entropy should be further developed to describe and quantify multipartite QE.