### v4

# Reduction-induced Variation of Partial Von Neumann Entropy

Abstract
The organization and structure of bipartite mixed-state quantum entanglement (QE) are more complex and less well understood compared to bipartite pure-state QE. Bipartite mixed-state QEs and their measures play a crucial role in both theory and practical applications. Some existing measures involve quantifying the minimum QE and reflect the inherently complex nature of their computation, while others are only applicable to highly limited-dimensional quantum systems. In this context, we propose a method termed Reduction-induced Variation of Partial Von Neumann Entropy to quantify QE in any bipartite states, particularly focusing on bipartite mixed states. Partial Von Neumann Entropy is merely a special case of this method, specifically designed to measure QE in bipartite pure states. This method exhibits minimal computational complexity and broad applicability. Its intuitive and clear physical representation, along with easy computation and wide applicability, facilitates exploring its potential applications. Furthermore, we present examples to demonstrate the superiorities of this method in identifying bipartite QE by comparing with other existing bipartite mixed-state QE measures through both their physical implications and mathematical structures.
Keywords:Quantum Entanglement (QE); Bipartite Mixed-state QE; Reduction-induced Variation of Partial Von Neumann Entropy
pacs:
03.65.Ud; 75.10.Pq; 05.30.-d
I Introduction
In 1935, “Spooky Action at a Distance”described as a strange phenomenon of quantum mechanics calls into question the completeness of the quantum theory by the famous physicists Einstein, Podolsky and Rosen (EPR) epr . And then in the same year, the concept of quantum entanglement (QE) was first proposed clearly by Schrödinger sd , through rigorous mathematical formalism, he demonstrated that particles in such entangled states cannot be individually described but must be treated as an integral composite system. Subsequently, in 1964, Bell realized that QE can produce experimentally verifiable deviations from classical physics in quantum mechanics bell . However, even to this day, its mystery still challenges our understanding of the laws of nature.
The type or degree of QEs are often directly related to their revolutionary applications including quantum cryptography kk ; ng , quantum teleportation ch ; rc ; nr , quantum superdense coding nr ; sj , measurement-based quantum computation rr ; fa and quantum enhanced sensing tc ; rcp ; lp . Therefore, a deeper understanding and accurate metric of QE have become important prerequisites for the development of quantum information science and quantum technology. However, to date, the quantum mechanical structure of QE in bipartite mixed states remains unclear; let alone the organization and structure of multipartite QEs. Here, our primary focus is on QEs in bipartite mixed states.
Nonseparable bipartite systems exhibit QE epr ; sd , which can be quantified using various measures. These include the Partial Von Neumann Entropy ch1 ; ch2 , applicable only to bipartite pure states, and Concurrence, which is confined to $2\times 2$ quantum systems wk ; au ; ka ; mf ; xjw . Additionally, Entanglement of Formation ch1 ; ch2 and Relative Entropy of Entanglement vv1 ; vv2 for bipartite quantum states, involve quantifying the minimum QE, and their assessments generally entail considerable computational burdens, especially for bipartite mixed-state QE. In general, QEs in bipartite pure states are more intuitive and easier to comprehend, whereas those in bipartite mixed states are more complicated and less well understood. The quantification of QEs in bipartite mixed states is more complex and challenging due to the contributions from classical probability distributions, making the definition and computation of QEs more intricate.
Mixed states are more reflective of real-world scenarios, as factors like environmental conditions, quantum measurements and thermal influences can lead quantum systems to exist in mixed states. Hence, the quantification of QE in bipartite mixed states is more meaningful in practice, aligning more effectively with experimental conditions and practical applications. Hence, in practical research and applications, bipartite mixed-state QE and its measures hold significant practical importance, in addition to the theoretical significance rooted in the underlying structure of quantum mechanics.
In this context, our research aims to explore new metrics for quantifying bipartite mixed-state QE. We propose a method termed Reduction-induced Variation of Partial Von Neumann Entropy to quantify QE in any bipartite states. Partial Von Neumann Entropy is merely a special case of this method, serving as a measure of bipartite pure-state QE. The computational complexity associated with this approach is considerably low, making it highly efficient. Moreover, it has a wide range of applicability. The intuitive and clear physical representation, along with simple computation and wide range of applications, facilitates the exploration of its specific potential applications. Subsequently, we provide examples to demonstrate the superiorities of this method in bipartite QE identification through a comparison with other established bipartite QE measures such as Partial Von Neumann Entropy, Relative Entropy of Entanglement, Entanglement of Formation and Concurrence considering both their physical implications and mathematical structures.
II existing four kinds of Bipartite QE measures
II.1 Partial Von Neumann Entropy for Bipartite Pure-state QE
For a pure bipartite state $\mid\Psi\rangle$, its density matrix $\rho=\mid\Psi\rangle\langle\Psi\mid$, a well-known and commonly used bipartite pure-state QE measure, the Partial Von Neumann Entropy (PVNE) ch1 ; ch2 is defined as
$\operatorname{PVNE}(\rho)=S(\rho_{i})=-\operatorname{Tr}\rho_{i}\log_{2}{\rho_{i}},$ (1)
where $\rho_{i}(i=1,2)$ represents its subsystem reduced density matrix. This Partial Von Neumann Entropy can effectively measure the degree of QE between bipartite pure states, while for bipartite mixed-state QE, it is ineffective.
II.2 Entanglement of Formation for Bipartite QE and Concurrence for two-qubit QE
For bipartite QE, especially for bipartite mixed-state QE, an extensively acknowledged type of measure, Entanglement of Formation (EOF) ch1 ; ch2 was introduced and is defined as
$\operatorname{EOF}(\rho)=\min_{\{p_{j},\mid\Psi_{j}\rangle\}}\sum_{j}p_{j}\operatorname{PVNE}(\mid\Psi_{j}\rangle\langle\Psi_{j}\mid),$ (2)
over all possible ensemble realizations of the bipartite density matrix $\rho=\sum_{j}p_{j}\mid\Psi_{j}\rangle\langle\Psi_{j}\mid$, where $p_{j}>0$ and $\sum_{j}p_{j}=1$. Entanglement of Formation represents the minimum entanglement resources required to construct this state and reflects that the involving computation is inherently complex. While for bipartite pure-state QE, Entanglement of Formation simplifies to the Partial Von Neumann Entropy
$\operatorname{EOF}(\mid\Psi\rangle)=\operatorname{PVNE}(\mid\Psi\rangle)=S(\rho_{i})=-\operatorname{Tr}\rho_{i}\log_{2}{\rho_{i}}.$ (3)
However, for mixed-state QE, in general, this measure heavily relies on the decompositions of pure states, with no general algorithm existing to determine the minimum decomposition to date. To address this issue and improve the Entanglement of Formation behavior in mixed states, Reference wk discovered that for two-qubit mixed states, the Entanglement of Formation reduces to
$\operatorname{EOF}(\rho)=H(z)=-z\log(z)-(1-z)\log(1-z),$ (4)
where the argument z is given by
$z=\frac{1+\sqrt{1-C^{2}}}{2},$ (5)
and $C(\rho)$ denotes Concurrence which is defined as wk ; au ; ka ; mf ; xjw
$\operatorname{C}(\rho)=\max\{0,\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4}\},$ (6)
where $\{\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\}$ represent the square roots of the eigenvalues of the matrix $\rho\tilde{\rho}$ arranged in decreasing order. The matrix $\tilde{\rho}$ is defined as
$\tilde{\rho}=(\sigma_{y}\otimes\sigma_{y})\rho^{*}(\sigma_{y}\otimes\sigma_{y}),$ (7)
where $\sigma_{y}$ is the Pauli matrix and $\rho^{*}$ denotes the complex conjugate of $\rho$. It is worth noting that, for two-qubit states, in itself, Concurrence is also a widely accepted and comparably practical measure of QE.
II.3 Relative Entropy of Entanglement for Bipartite QE
Another well-known and commonly used bipartite QE measure, especially for bipartite mixed-state QE, Relative Entropy of Entanglement (REE) vv1 ; vv2 is defined as
$\operatorname{REE}=S(\rho|\sigma)=\min_{\sigma\in\mathcal{S}}[-\operatorname{Tr}\rho\log_{2}{\sigma}-S(\rho)],$ (8)
where $\mathcal{S}$ represents a set of separable bipartite density matrix. On one hand, Relative Entropy of Entanglement measures the shortest “distance”from a given quantum entangled state to the set of separable states, which is the geometric interpretation of this method and reflects that the involving computation is inherently complex. On the other hand, while for bipartite pure QE states, the Relative Entropy of Entanglement simplifies to the Partial Von Neumann Entropy
$\operatorname{REE}$ $=S(\rho|\sigma)=\min_{\sigma\in\mathcal{S}}[-\operatorname{Tr}\rho\log_{2}{\sigma}]$ (9)
$=\operatorname{PVNE}(\rho)=S(\rho_{i})=-\operatorname{Tr}\rho_{i}\log_{2}{\rho_{i}}.$
The four kinds of bipartite QE measures mentioned above exhibit various flaws and limitations. For example, Partial Von Neumann Entropy is limited to pure bipartite quantum systems; Entanglement of Formation and Relative Entropy of Entanglement involve determining the minimum QE required for a given state, reflecting the inherent complexity of their computation. While Concurrence is applicable only to $2\times 2$ quantum systems. The organization and structure of bipartite mixed-state QE are more complicated and less well understood than those of pure-state QE. Quantifying QE in mixed states poses additional challenges due to the contributions from classical probability distributions, which complicate both the definition and computation of QE measures. In this context, our research aims to explore new measures for bipartite mixed-state QE in detail, as outlined below.
III Reduction-induced Variation of Partial Von Neumann Entropy
In any bipartite QE system, all the information is contained within its wave function or density matrix. When a subsystem is obtained by partial tracing or averaging out the freedom of the other subsystem, some information is inevitably lost. Therefore, the concept of Reduction-induced Variation of Partial Von Neumann Entropy is presented to quantify bipartite QE, which holds specific importance and practical significance. In the bipartite density matrix $\rho$, the density matrix of the $i$-th $(i=1,2)$ party, only preserves all the information of the $i$-th party from $\rho$; this bipartite $i$-th part density matrix is denoted as $\rho(i)$ and differs from its reduced density matrix $\rho_{i}$. To elaborate further, let’s begin with bipartite qubit systems. As is well known, the Bell entangled states are
$\Psi_{\pm}=\frac{\sqrt{2}}{2}|11\rangle\pm\frac{\sqrt{2}}{2}|00\rangle,$ (10)
and
$\Phi_{\pm}=\frac{\sqrt{2}}{2}|10\rangle\pm\frac{\sqrt{2}}{2}|01\rangle.$ (11)
And their corresponding density matrices are respectively
$\rho=\left[\begin{array}[]{cccc}\frac{1}{2}&0&0&\frac{1}{2}\\ 0&0&0&0\\ 0&0&0&0\\ \frac{1}{2}&0&0&\frac{1}{2}\\ \end{array}\right]$ (12)
with the corresponding bipartite single-site density matrix and the single-site reduced density matrix provided as follows:
$\rho(i)=\left[\begin{array}[]{cc}\frac{1}{2}&\frac{1}{2}\\ \frac{1}{2}&\frac{1}{2}\\ \end{array}\right]$ (13)
and
$\rho_{i}=\left[\begin{array}[]{cc}\frac{1}{2}&0\\ 0&\frac{1}{2}\\ \end{array}\right];$ (14)
and
$\rho=\left[\begin{array}[]{cccc}0&0&0&0\\ 0&\frac{1}{2}&\frac{1}{2}&0\\ 0&\frac{1}{2}&\frac{1}{2}&0\\ 0&0&0&0\\ \end{array}\right]$ (15)
with the corresponding bipartite single-site density matrix and the single-site reduced density matrix specified as follows:
$\rho(i)=\left[\begin{array}[]{cc}\frac{1}{2}&\frac{1}{2}\\ \frac{1}{2}&\frac{1}{2}\\ \end{array}\right]$ (16)
and
$\rho_{i}=\left[\begin{array}[]{cc}\frac{1}{2}&0\\ 0&\frac{1}{2}\\ \end{array}\right].$ (17)
For the Bell entangled states, the Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE) is expressed as:
$\operatorname{RIVPVNE}(\rho)=S(\rho_{i})-S(\rho(i))=S(\rho_{i}).$ (18)
Inspired by this, we consider a general entangled bipartite qubit system. The density matrix that best reflects the QE of the bipartite qubit must satisfy the following form,
$\rho_{\operatorname{eff}}=\left[\begin{array}[]{cccc}\rho_{11,11}&0&0&\rho_{11,00}\\ 0&\rho_{10,10}&\rho_{10,01}&0\\ 0&\rho_{01,10}&\rho_{01,01}&0\\ \rho_{00,11}&0&0&\rho_{00,00}\\ \end{array}\right],$ (19)
that is to say
$\rho_{11,10}=\rho_{10,11}=\rho_{11,01}=\rho_{01,11}$
$=$ $\rho_{00,10}=\rho_{10,00}=\rho_{00,01}=\rho_{01,00}=0,$ (20)
which serves as a necessary condition. The system’s density matrix can take different forms based on different single-site orthonormal basis vectors $\{\cos\theta|1\rangle-\sin\theta|0\rangle,\cos\theta|0\rangle+\sin\theta|1\rangle,0\leqslant\theta<2\pi\}$}. Therefore, we first select an appropriate orthonormal basis such that the density matrix is expressed in a form that most effectively represents the entangled state $\rho_{\operatorname{eff}}$. Under these conditions, the entropy of the bipartite $i$-th part density matrix reaches its maximum value $S_{\operatorname{Max}}(\rho(i)_{\operatorname{eff}})$ depending on the choice of single-site orthonormal basis vectors. The same approach is also applied to bipartite high-dimensional quantum systems. The Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE) of one-sided density matrix is defined as follows:
$\operatorname{RIVPVNE}(\rho)=S(\rho_{i})-S_{\operatorname{Max}}(\rho(i)_{\operatorname{eff}}),$ (21)
This measure vanishes if and only if the state is separable, and on average, does not experience an increase under Stochastic Local Operations and Classical Communications, thereby establishing itself as an entanglement monotone G.V ; Y.C . Consequently, it can be employed to quantify the QE degree in bipartite systems. Notably, the computational complexity associated with this method is minimal, making it highly efficient. Moreover, it has a wide range of applicability. The intuitive and clear physical picture, combined with simple computation and wide range of applications, facilitates the exploration of its specific potential applications. For any bipartite pure-state QE, this measure reduces to the Partial Von Neumann Entropy
$\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=-\operatorname{Tr}\rho_{i}\log_{2}{\rho_{i}}.$ (22)
Partial Von Neumann Entropy is merely a special case of this method, serving as a measure of bipartite pure-state QE. Due to the locality of this measure, in general only for symmetric states, it has symmetry about different partition. Subsequently, we provide examples to demonstrate the superiorities of this method in identifying QE through a comparison with other existing bipartite QE measures such as Partial Von Neumann Entropy, Relative Entropy of Entanglement, Entanglement of Formation and Concurrence taking into account both their physical implications and mathematical structures, as detailed below.
IV Its Superiorities in comparison to existing bipartite QE measures
We initially consider the symmetric pure Bell-like states
$|\Psi\rangle=\cos\alpha|11\rangle\pm\sin\alpha|00\rangle$ (23)
and
$|\Phi\rangle=\cos\alpha|10\rangle\pm\sin\alpha|01\rangle$ (24)
with $0\leqslant\alpha\leqslant\frac{\pi}{2}$. Consequently, the Reduction-induced Variation of Partial Von Neumann Entropy, as well as the Entanglement of Formation and Relative Entropy of Entanglement, simplifies to the Partial Von Neumann Entropy:
$\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)$
$=-(\cos^{2}\alpha\log_{2}{\cos^{2}\alpha}+\sin^{2}\alpha\log_{2}{\sin^{2}\alpha}).$ (25)
Additionally, Concurrence and Negativity can be expressed as
$\operatorname{C}(\rho)=|\sin{2\alpha|}.$ (26)
This highlights the interference properties and indicates a clear and intuitive relationship between QE and quantum coherence. The previously mentioned measures or witnesses are illustrated in Fig.1. For the specified pure states, it is observed that when $\alpha=\frac{\pi}{4}$, all the measures satisfy: $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=1$; whereas for $\alpha=0,\frac{\pi}{2}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0$; and for the ranges $0<\alpha<\frac{\pi}{2}$ and $\frac{\pi}{4}<\alpha<\frac{\pi}{2}$, the relationships $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)<\operatorname{C}(\rho)$ hold. Clearly, the monotonicities of these measures are fully consistent. The Reduction-induced Variation of Partial Von Neumann Entropy, Entanglement of Formation and Relative Entropy of Entanglement all reduce to the Partial Von Neumann Entropy, but is less than or equal to Concurrence which here signifies quantum coherence, because of differences in their physical implications and mathematical structures.
Fig. 1: Reduction-induced Variation of Partial Von Neumann Entropy as well as other QE measures such as the Partial Von Neumann Entropy, Entanglement of Formation, Relative Entropy of Entanglement and Concurrence varying with $\alpha$. When $\alpha=\frac{\pi}{4}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=1$; whereas for the values $\alpha=0,\frac{\pi}{2}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0$; and in the ranges $0<\alpha<\frac{\pi}{4}$ and $\frac{\pi}{4}<\alpha<\frac{\pi}{2}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{PVNE}(\rho)=S(\rho_{i})=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)<\operatorname{C}(\rho)$.
We next consider the symmetric bipartite states
$\rho=\frac{\sin^{2}\alpha}{2}\mid 10\pm 01\rangle\langle 10\pm 01|+\cos^{2}\alpha|00\rangle\langle 00\mid$ (27)
with $0\leqslant\alpha\leqslant\pi$. The first term corresponds to a pure-state QE with a probability of $\sin^{2}\alpha$, while the second term represents a non-entangled pure state with a probability of $\cos^{2}\alpha$. Hence, the bipartite mixed-state QE is less than $\sin^{2}\alpha$. Next, the Reduction-induced Variation of Partial Von Neumann Entropy (RIVPVNE) is computed as
$\operatorname{RIVPVNE}=S(\rho_{i})-S_{\operatorname{Max}}(\rho(i)_{\operatorname{eff}})$
$=-(\frac{3+\cos{2\alpha}}{4}\log_{2}{\frac{3+\cos{2\alpha}}{4}}+\frac{\sin^{2}\alpha}{2}\log_{2}{\frac{\sin^{2}\alpha}{2}})$
$+\frac{2-\sqrt{3+\cos{4\alpha}}}{4}\log_{2}{\frac{2-\sqrt{3+\cos{4\alpha}}}{4}}$
$+\frac{2+\sqrt{3+\cos{4\alpha}}}{4}\log_{2}{\frac{2+\sqrt{3+\cos{4\alpha}}}{4}},$ (28)
in addition to other QE measures like the Entanglement of Formation given by
$\operatorname{EOF}(\rho)=$ $-\frac{1-\sqrt{1-\sin^{4}\alpha}}{2}\log_{2}{\frac{1-\sqrt{1-\sin^{4}\alpha}}{2}}$ (29)
$-\frac{1+\sqrt{1-\sin^{4}\alpha}}{2}\log_{2}{\frac{1+\sqrt{1-\sin^{4}\alpha}}{2}},$
and the Relative Entropy of Entanglement as well as Concurrence given by
$\operatorname{REE}(\rho)=\operatorname{C}(\rho)=\sin^{2}\alpha,$ (30)
which represents the upper limit of the bipartite mixed-state QE as derived from the mathematical structure of the system’s density matrix. The aforementioned four kinds of bipartite QE measures are depicted in Fig.2. Clearly, for the specified bipartite quantum systems, the monotonic behavior of these measures is consistent. Notably, when $\alpha=\frac{\pi}{2}$, the following relations hold: $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=1$; whereas for $\alpha=0,\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0$. While within the ranges $0<\alpha<\frac{\pi}{2}$ and $\frac{\pi}{2}<\alpha<\pi$, the relations become $\operatorname{RIVPVNE}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$. The Relative Entropy of Entanglement and Concurrence, representing the upper limit of QE, are the largest among these measures, and reflect that they may not accurately reflect the degree of bipartite mixed-state QE. Conversely, the Reduction-induced Variation of Partial Von Neumann Entropy is the smallest, with the Entanglement of Formation falling in between. These differences in metric values, arise from their distinct physical implications and mathematical structures. Regarding the bipartite QE degree metric for mixed states, it is concluded that the Reduction-induced Variation of Partial Von Neumann Entropy and Entanglement of Formation exhibit advantages over the Relative Entropy of Entanglement and Concurrence. When considering bipartite QE degree metric in mixed states, the computational complexity and range of applicability, the Reduction-induced Variation of Partial Von Neumann Entropy is superior to the Entanglement of Formation, Relative Entropy of Entanglement and Concurrence.
Fig. 2: Reduction-induced Variation of Partial Von Neumann Entropy as well as other QE measures such as Entanglement of Formation, Relative Entropy of Entanglement and Concurrence varying with $\alpha$. When $\alpha=\frac{\pi}{2}$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=1$; whereas for $\alpha=0,\pi$, $\operatorname{RIVPVNE}(\rho)=\operatorname{EOF}(\rho)=\operatorname{REE}(\rho)=\operatorname{C}(\rho)=0$. While within the ranges $0<\alpha<\frac{\pi}{2}$ and $\frac{\pi}{2}<\alpha<\pi$, $\operatorname{RIVPVNE}(\rho)<\operatorname{EOF}(\rho)<\operatorname{REE}(\rho)=\operatorname{C}(\rho)$.
V Conclusion
In this paper, we propose a method called Reduction-induced Variation of Partial Von Neumann Entropy to quantify QE in any bipartite sates, especially for mixed states. Partial Von Neumann Entropy is merely a special case of this method, serving as a measure of bipartite pure-state QE. This method exhibits minimal computational complexity, making it highly efficient. Moreover, it has a wide range of applicability. Its intuitive and clear physical representation, along with simple computation and wide range of applications, indicates the ease of exploring its specific potential applications. By considering the degree of bipartite QE, along with the computational complexity and range of applicability, the Reduction-induced Variation of Partial Von Neumann Entropy proves to be a good measure for assessing bipartite mixed-state QE compared to other established methods like Entanglement of Formation, Relative Entropy of Entanglement and Concurrence. In future studies, the conceptual framework and line of thought behind the Reduction-Induced Variation of Partial Von Neumann Entropy deserves to be further developed to describe and quantify multipartite QE.